OpenStudy (anonymous):

How much force, in Newtons, is applied to a 4 kg bowling ball that goes from 0 m/s to 5m/s in .5 seconds

OpenStudy (mathmate):

hints: F=ma, where F=force in newtons, m=mass in kg, acceleration=\((v_f-v_i)/\Delta t\) = (5-0)/0.5 m/s^2

OpenStudy (anonymous):

would it be 20 newtons?

OpenStudy (mathmate):

Remember to divide the change in velocity by the time lapse: F=ma = 4 ( (5-0)/0.5 ) = ?